# Diffusion Feedback Helps CLIP See Better

### v1

@@ 3.3 Diffusion’s Condition Design
Visual Recap Density. Although introducing the visual dense recap scheme appears straightforward, the density of the recap is crucial. If the density is too high (i.e., introducing too many local tokens), the richness of the condition information approaches its maximum, significantly reducing the difficulty of the entire reconstruction task. This results in CLIP’s representation requiring minimal optimization to easily complete the reconstruction, thereby greatly limiting the upper bound of CLIP’s optimized capabilities. Conversely, if the visual recap density is too low (i.e., retaining only the class token or introducing too few local tokens), CLIP’s optimization process will struggle with the high difficulty of the reconstruction task, failing to adequately learn the expected detailed visual representations. This intuitive point is confirmed in the ablation experiments in Sec. 4.5. Specifically, ensuring the visual class token is always present in the condition, we introduce randomly selected local token features with approximately 15% and 30% probabilities for OpenAI CLIP (Radford et al., 2021b) at 224 and 336 image resolutions, respectively. For the SigLIP ViT-SO-14 (Zhai et al., 2023) at 224 and 384 image sizes, we incorporate local token features obtained through average pooling with local window sizes of 1$\times$6 and 1$\times$10, respectively. For the remaining two baselines (Fang et al., 2023; Xu et al., 2023a), we include all local token features as part of the condition. Apart from DFN (Fang et al., 2023) only using visual class token, all other models incorporate local features consistent with the training stage conditions during inference, combining them with class token to fully leverage the detailed representations captured by the enhanced CLIP.

### v2

@@ 3.3 Diffusion’s Condition Design
Visual Recap Density. Although the visual dense recap scheme appears straightforward, the density of the recap is crucial. If the density is too high (i.e., introducing too many local tokens), the richness of the condition information approaches its maximum, greatly reducing the difficulty of reconstruction task. This results in CLIP’s representation requiring minimal optimization to easily complete the reconstruction, limiting the upper bound of CLIP’s optimized capabilities. Conversely, if the recap density is too low (i.e., retaining only class token or introducing few local tokens), CLIP’s optimization process will struggle with the high difficulty of reconstruction, failing to adequately learn the expected detailed visual representations. This intuitive point is confirmed in Sec. 4.5. Specifically, ensuring the visual class token is always present in the condition, we introduce randomly selected local token features with approximately 15% and 30% probabilities for OpenAI CLIP (Radford et al., 2021b) at 224 and 336 resolutions. For the SigLIP ViT-SO-14 (Zhai et al., 2023) at 224 and 384 image sizes, we incorporate local token features obtained through average pooling with local window sizes of 1$\times$6 and 1$\times$10, respectively. Except that introducing 50% randomly selected patch tokens into the condition for DFN ViT-H-14/336, for the remaining baselines (Fang et al., 2023; Xu et al., 2023a), we include all local token features as part of the condition. Apart from DFN ViT-H-14/224 and SigLIP ViT-SO-14/224&384 (Zhai et al., 2023) only using visual class token, all other models incorporate local features consistent with the training stage conditions during inference, combining them with class token to fully leverage the detailed representations captured by the enhanced CLIP.